OpenStudy (anonymous):

make invites 14 guests .he wants every guest to eat 3\4 of ib turkey .how many kg of turkey he needs(remember 1ib=450g) plz help me .thank you

OpenStudy (dan815):

1 lb not ib lol

OpenStudy (anonymous):

oh yes.Imade a mistake.thank you

OpenStudy (anwaarullah):

14*3/4=10.5 lb convert to grams 10.5*450=4725 grams now to Kg 4725/1000= 4.725 kg

OpenStudy (anonymous):

thanks pruv